Researchers measured the seated systolic blood pressure of 47 healthy white males. The resulting publication reports x ̅= 114.9 mm Hg and s = 9.3 mm Hg. Give a 95% confidence interval for the mean blood pressure in the population from which the subjects were recruited. Question 7 options: A) (105.6, 124.2) B) (112.2, 117.6) C) (111.4, 118.4) D) (112.7, 117.1) E) (114.4, 115.4)
Answer #1

Solution :

Given that,

= 114.9

s =9.3

n =47

Degrees of freedom = df = n - 1 =47 - 1 = 46

a ) At 95% confidence level the t is ,

= 1 - 95% = 1 - 0.95 = 0.05

  / 2 = 0.05 / 2 = 0.025

t /2,df = t0.025,46 = 2.013

Margin of error = E = t/2,df * (s /n)

= 2.013* ( 9.3/ 47)

=2.7

The 95% confidence interval estimate of the population mean is,

- E < < + E

114.9 -2.7 < < 114.9+ 2.7

112.2 < < 117.6

(112.2 , 117.6 )